White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates:

Department
Cutting Finishing
Direct labor-hours 7,900 86,000
Machine-hours 64,600 2,600
Total fixed manufacturing overhead cost $ 390,000 $ 410,000
Variable manufacturing overhead per machine-hour $ 3.00
Variable manufacturing overhead per direct labor-hour $ 3.75

Required:

1. Compute the predetermined overhead rate for each department.

2. The job cost sheet for Job 203, which was started and completed during the year, showed the following:

Department
Cutting Finishing
Direct labor-hours 4 18
Machine-hours 84 5
Direct materials $ 790 $ 370
Direct labor cost $ 80 $ 360

3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates?

1. Compute the predetermined overhead rate for each department. (Round your answers to 2 decimal places.)

2. Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203. (Round intermediate calculations to 2 decimal places and final answer to the nearest dollar amount.)

3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates?

Yes
No
Answer #1

SOLUTION

1. Cutting Department

Predetermined overhead rate = [Total fixed manufacturing overhead costs / Machine hours] + Variable manufacturing overhead per machine-hour

= [390,000 / 64,600] + $3

= 6.04 + 3 = 9.04 per machine hour

Finishing department

Predetermined overhead rate = [Total fixed manufacturing overhead costs / Direct labor hours] + Variable manufacturing overhead per machine-hour

= [410,000 / 86,000] + $3.75

= 4.77 + 3.75 = 8.52 per direct labor hour

2.

Cutting department Finishing department
Direct materials 790 370
Direct labor cost 80 360
Manufacturing Overhead [(84*$9.04), (18*$8.52)] 759.36 153.36
Total Manufacturing cost 1,629.36 883.36

3. Yes, if plant wide overhead rate would be used allocation of overhead would be changed since in present case only respective hours in particular department is used, and for plant wide rate per direct labor hour would be used that will be lower than present case.